A 23 gram bullet is fired from a gun with a speed of 225m/s. If the gun has a mass of 0.7kg. What is the recoil speed of the gun?


1
Expert's answer
2021-04-26T20:15:56-0400

According to the momentum conservation law, the momentum of the gun is equal to the momentum of the bullet (by module):


"m_bv_b = m_gv_g"

where "m_b = 23g = 0.023kg, m_g = 0.7kg" are the masses of the bullet and the gun respectively, and "v_b = 225m\/s, v_g" are the velocities of the bullet and the gun respectively. Expressing "v_g", obtain:


"v_g = \dfrac{m_bv_b}{m_g}\\
v_g = \dfrac{0.023kg\cdot 225m/s}{0.7kg} \approx 7.4 m/s"

Answer. 7.4 m/s.